# On putative self-similarity for incompressible 3D Euler

### v1

@@ 4.4. Failure of the outgoing property: all fixed points on the axis of symmetry
From Theorem 4.3 we know that if a sufficiently smooth [^12] By this smoothness we mean: $U$ is globally $C^{2}$ smooth, and $\Omega$ is real-analytic in the neighborhood of all (isolated) fixed points $(r,z)$ of the meridional-plane ODE (4.4). axisymmetric similarity profile $U\not\equiv 0$ solving (4.1) were to exist for some $\gamma<1/2$, then the outgoing property (4.6) must fail.
The most natural case is when the failure of the outgoing property results in the existence of fixed points points of the ODE system (4.4), other than the origin. [^13] For instance, the outgoing property (4.6) can fail at a point $(0,z)\neq(0,0)$ on the axis of symmetry if and only if $(0,z)$ solves (4.7), i.e., $(0,z)$ is a fixed point for (4.4). Due to Theorem 4.4 the swirl velocity must vanish at all these fixed points. By Remark 4.2, this vanishing of the swirl velocity would be dynamically stable for instance if all fixed points of the ODE system (4.4) would lie on the axis of symmetry $\{r=0\}$. [^14] For 3D incompressible Navier-Stokes, any putative finite time axisymmetric singularity must occur on the axis of symmetry [6]. Hence, if we are to lift axisymmetric self-similar blowups from incompressible 3D Euler to incompressible 3D Navier-Stokes, it is natural to search for Euler singularities which lie on $\vec{e}_{z}$.
In this section we analyze putative self-similar solutions of 3D incompressible Euler in axisymmetry (4.1), for which the outgoing property fails, and for which all solutions of (4.7) lie on $\vec{e}_{3}$.
@@
Let $U\not\equiv 0$ be a $C^{2}$ smooth axisymmetric globally self-similar velocity profile for 3D incompressible Euler equations, solving (4.1), satisfying (3.5) and (3.8). Assume that all fixed points of the meridional Lagrangian flow (4.4) are isolated, and lie on the axis of symmetry $\{r=0\}$. Assume that $\Omega$ is real-analytic in a neighborhood of each such fixed point. Then $\gamma\geq 1/2$.
@@
The fixed points of (4.4) on the axis of symmetry $\{r=0\}$ are precisely the zeros of the function
$h(z):=\gamma z+U_{z}(0,z).$
@@
Because $U(0)=0$ (cf. (3.5)), the origin is always a zero: $h(0)=0$. Furthermore, the sublinearity assumption (3.5) implies that $h(z)\sim\gamma z$ as $|z|\to\infty$, so $h$ has at most finitely many zeros (they are isolated). Throughout this proof, we denote
$g(z):=\partial_{z}U_{z}(0,z).$
Since $U_{r}(0,z)=0$, we have that $\lim_{r\to 0^{+}}\frac{1}{r}U_{r}(r,z)=\partial_{r}U_{r}(0,z)$, and thus incompressibility (the last component of (4.1)) gives $\partial_{r}U_{r}(0,z)=-\frac{1}{2}\partial_{z}U_{z}(0,z)=-\frac{1}{2}g(z)$; this fact will be used freely in the proof. We also note also that the definitions of $g$ and $h$ imply
$h^{\prime}(z)=\gamma+g(z).$
We denote
$\phi(z):=\Omega_{z}(0,z).$
Note that second equation in (4.3) restricted to $r=0$ implies that
$\phi(1-g)+h\phi^{\prime}=0,$ (4.8)
pointwise at all $z\in\mathbb{R}$.
Assume by contradiction that $\gamma<1/2$. We will prove $\Omega\equiv 0$, contradicting the blowup scenario.
Step 1: Ruling out $g=1$ at all axis fixed points. Denote the fixed point at the origin by $z_{0}=0$, and label the zeros of $h(z)$ in increasing order as $z_{-M}<\ldots z_{-1}<0<z_{1}<\ldots<z_{N}$, for some $N,M\in\mathbb{N}_{0}$. We claim that $g(z_{j})=\partial_{z}U_{z}(0,z_{j})\neq 1$ for all $M\leq j\leq N$.
Suppose $g(z_{j})=1$ for some $j$. On the axis of symmetry $U_{r}\equiv 0$ and $\Omega_{\theta}\equiv 0$, which implies $\partial_{z}U_{r}\equiv 0$ and $\partial_{r}U_{z}\equiv 0$. Therefore, the linearization of (4.4) at $(0,z_{j})$ is diagonal, and with $\partial_{r}U_{r}(0,z)=-\frac{1}{2}g(z)$ we deduce
$\gamma\mathrm{Id}+\begin{pmatrix}\partial_{r}U_{r}&\partial_{z}U_{r}\\ \partial_{r}U_{z}&\partial_{z}U_{z}\end{pmatrix}(0,z_{j})=\begin{pmatrix}\gamma-\frac{g(z_{j})}{2}&0\\ 0&\gamma+g(z_{j})\end{pmatrix}=\begin{pmatrix}\gamma-\frac{1}{2}&0\\ 0&\gamma+1\end{pmatrix}$.
Because we have assumed $\gamma<1/2$, the eigenvalue of the linearization at $(0,z_{j})$ corresponding to the radial direction is stable ($\gamma-\frac{1}{2}<0$), while the eigenvalue corresponding to the axial direction is unstable ($\gamma+1>0$). Thus $(0,z_{j})$ is a hyperbolic saddle for the meridional ODE flow (4.4) in the half-plane $\{r\geq 0,z\in\mathbb{R}\}$. The one-dimensional stable manifold $\mathcal{W}^{s}$ of this fixed point extends into the open domain $\{r>0,z\in\mathbb{R}\}$.
Next, we trace $\mathcal{W}^{s}$ backward in time $\tau$. Due to (3.32), for $\gamma<1/2$ the Bernoulli function $\mathcal{H}$ defined in (3.30) is a strict Lyapunov function: it is strictly decreasing along non-stationary forward trajectories, respectively strictly increasing along non-constant backwards trajectories lying in $\{r>0,z\in\mathbb{R}\}$. Therefore:
  • There are no periodic orbits for the meridional flow (4.4). Indeed, on a periodic orbit $(\gamma r+U_{r},\gamma z+U_{z})$ is nonzero, so $\mathcal{H}$ would strictly decrease over one period, contradicting periodicity.
  • The backward trajectory on $\mathcal{W}^{s}$ (emanating from $(0,z_{j})$ at time $\tau=+\infty$) remains bounded as $\tau\to$-$\infty$. On the one hand, along backward trajectories $\mathcal{H}$ increases from its value at the fixed point, $\mathcal{H}(0,z_{j})$, cf. (3.32). On the other hand, the far-field asymptotic in (3.33) gives $\mathcal{H}(y)\to$-$\infty$ as $|y|=(r^{2}+z^{2})^{1/2}\to\infty$. Hence the backward trajectory on $\mathcal{W}^{s}$ cannot escape to infinity.
By the Poincaré–Bendixson theorem for the two-dimensional meridional flow (restricted to the invariant closed half-plane $\{r\geq 0,z\in\mathbb{R}\}$), the $\alpha$-limit set of the bounded backward trajectory $\mathcal{W}^{s}$ is a compact, connected, flow-invariant set. As periodic orbits are excluded, and all fixed points lie on the axis, the $\alpha$-limit is an axis fixed point $(0,z_{k})$ for some $M\leq k\leq N$. We have thus obtained either a heteroclinic or a homoclinic orbit through the set $\{r>0,z\in\mathbb{R}\}$ connecting axis fixed points. This contradicts Corollary 4.8, and therefore, $g(z_{j})\neq 1$ for every axis fixed point $z_{j}$.
Step 2: $\phi(z)=0$ for all $z\in\mathbb{R}$. Recall that $\phi(z):=\Omega_{z}(0,z)$ obeys a first-order linear ODE (4.8), with singularities at the zeros of $h$.
Consider the largest positive zero $z_{N}$ of $h$; if no positive zero other than $z_{0}=0$ exists, the arguments below apply at the origin. Since $h(z)\sim\gamma z>0$ for $z\gg 1$ and $h(z_{N})=0$, maximality implies $h(z)>0$ for all $z>z_{N}$, and thus $h^{\prime}(z_{N})\geq 0$. Equivalently, $g(z_{N})\geq-\gamma$. From Step 1 we know that $g(z_{N})\neq 1$, so evaluating (4.8) at $z=z_{N}$ gives $\phi(z_{N})=0$. We claim that $\phi^{(n)}(z_{N})=0$ for all $n\geq 0$. We proceed by induction on $n$, the base case $n=0$ being established. Assume $\phi^{(k)}(z_{N})=0$ for all $k\leq n-1$, with $n\geq 1$. Applying $\partial_{z}^{n}$ to (4.8), evaluating at $z=z_{N}$, using that $h(z_{N})=0$, and using that by induction all lower-order derivatives of $\phi$ vanish at $z_{N}$ (akin to the proof of Proposition 3.8), we arrive at
$\phi^{(n)}(z_{N})\bigl(1+n\gamma+(n-1)g(z_{N})\bigr)=0.$ (4.9)
The coefficient of $\phi^{(1)}(z_{N})$ is $1+\gamma>0$, and thus $\phi^{(1)}(z_{N})=0$. For $n\geq 2$, because $g(z_{N})\geq-\gamma$, the coefficient of $\phi^{(n)}(z_{N})$ is bounded from below by $1+n\gamma-(n-1)\gamma=1+\gamma>0$. We similarly deduce from (4.9) that $\phi^{(n)}(z_{N})=0$, completing the induction.
Since $U$ (and hence $\phi$) is real-analytic in a neighborhood of $z_{N}$ by assumption, we deduce that $\phi\equiv 0$ on an interval $(z_{N}-\epsilon_{N},z_{N}+\epsilon_{N})$ for some $\epsilon_{N}>0$. On the interval $(z_{N},\infty)$, we have $h(z)>0$ (no further zeros), so (4.8) is a regular first-order linear ODE. Since $\phi\equiv 0$ on $(z_{N},z_{N}+\epsilon_{N})$, uniqueness of ODE solutions implies $\phi\equiv 0$ on $(z_{N},\infty)$. Similarly, on the open interval $(z_{N-1},z_{N})$, the function $h$ does not vanish, so (4.8) is regular. As $\phi\equiv 0$ on $(z_{N}-\epsilon_{N},z_{N})\subset(z_{N-1},z_{N})$, uniqueness of ODE solutions implies $\phi\equiv 0$ also on the open interval $(z_{N-1},z_{N})$. By continuity of $\phi$, we have that $\phi(z_{N-1})=0$ and $\phi^{(n)}(z_{N-1}^{+})=0$ for all $n\geq 0$. Because $\phi$ is real-analytic at $z_{N-1}$ by assumption, the vanishing of all (right) derivatives forces $\phi\equiv 0$ on a full neighborhood $(z_{N-1}-\epsilon_{N-1},z_{N-1}+\epsilon_{N-1})$. We repeat this procedure: the vanishing of $\phi$ is pass from the neighborhood of a fixed point $z_{i}$ to the interval $(z_{i-1},z_{i})$ by ODE uniqueness, then the vanishing property is extended across the singular point $z_{i}$ by analyticity. After finitely many steps, we reach $z_{-M}$. Since $h\sim\gamma z>0$ for $z\ll-1$ and $h(z_{-M})=0$, the minimality of $z_{-M}$ implies that $h(z)<0$ for all $z<z_{-M}$. Via (4.8) we finally prove tat $\phi$ vanishes identically on $(-\infty,z_{-M})$, thereby concluding the proof of
$\phi(z)=\Omega_{z}(0,z)=0,\qquad\mbox{for all }z\in\mathbb{R}.$
Step 3: Every backward trajectory originating in $\{r>0,z\in\mathbb{R}\}$ approaches an axis fixed point. Fix an arbitrary point $(r_{0},z_{0})$ with $r_{0}>0$, and consider its backward trajectory $\Gamma_{0}=\{(R(\tau),Z(\tau))\colon\tau\leq 0\}=\{(R(r_{0},z_{0},\tau),Z(r_{0},z_{0}\tau))\colon\tau\leq 0\}$, as $\tau\to$-$\infty$.
@@
We first claim that $\Gamma_{0}$ is contained in a bounded set. Because $\mathcal{H}$ is non-increasing along forward trajectories (cf. (3.32)), backward in time $\mathcal{H}$ is non-decreasing. The far-field asymptotics (3.33) imply that the sublevel set $\mathcal{S}_{c}:=\{y\in\mathbb{R}^{3}:\mathcal{H}(y)\geq c\}$ is bounded for every $c\in\mathbb{R}$, since $\mathcal{H}(y)\to$-$\infty$ as $|y|\to\infty$ when $\gamma<1/2$. Setting $c=\mathcal{H}(r_{0},z_{0})$, the backward trajectory satisfies $\mathcal{H}(R(\tau),Z(\tau))\geq c$ for all $\tau\leq 0$, and is thus contained in $\Gamma_{0}\subset\mathcal{S}_{c}$.
@@
Next, we claim that the $\alpha$-limit set of $\Gamma_{0}$ is a single axis fixed point. The backward trajectory $\Gamma_{0}$ is bounded and lies in the two-dimensional meridional half-plane. By the Poincaré–Bendixson theorem, the $\alpha$-limit set $A$ is a nonempty, compact, connected, flow-invariant subset of the half-plane $\{r\geq 0,z\in\mathbb{R}\}$. As periodic orbits are excluded (see Step 1), $A$ must either be a single fixed point, or consist of multiple fixed points connected by heteroclinic/homoclinic orbits. Since all fixed points lie on the axis of symmetry $\{r=0\}$, any heteroclinic/homoclinic orbit in $A$ connecting axis fixed points would pass through $\{r>0\}$, violating Corollary 4.8. Therefore, $A=\{(0,z_{*})\}$ for some axis fixed point $z_{*}$, and the full trajectory converges: $(R(\tau),Z(\tau))\to(0,z_{*})$ as $\tau\to$-$\infty$.
Step 4: $\Omega\equiv 0$ in $\{r>0,z\in\mathbb{R}\}$. Fix an arbitrary point $(r_{0},z_{0})$ with $r_{0}>0$, and let $\Gamma_{0}$ be the backwards in $\tau$ trajectory from Step 3, with $(R(\tau),Z(\tau))\to(0,z_{*})$ as $\tau\to$-$\infty$. Since $\Omega_{r},\Omega_{\theta}$ (due to axisymmetry) and $\Omega_{z}$ (due to Step 2) all vanish on the axis of symmetry $\{r=0\}$, and since $\Omega$ is real-analytic at $(0,z_{*})$, there exists $\epsilon_{*}>0$ such that $\Omega$ vanishes identically in $\{(r,z)\colon r^{2}+(z-z_{*})^{2}<\epsilon_{*}^{2}\}$. Also, there exists $\tau_{*}>0$ such that $R(\tau)^{2}+(Z(\tau)-z_{*})^{2}<\epsilon_{*}^{2}$ for all $\tau>\tau_{*}$. Note that (4.4)–(4.5) is autonomous, so the backwards trajectory $\Gamma_{0}$ may be identified with the projection on the meridional plane of the back-to-labels map $Y^{-1}$ originating at $(r_{0},z_{0})$. As in the proof of Theorem 3.9, by appealing to the Cauchy formula (3.22) (in axisymmetric form) we write $\Omega(r_{0},z_{0})=M\Omega(R(\tau),Z(\tau))$ for a suitable stretching matrix $M$, and deduce $\Omega(r_{0},z_{0})=0$. ∎

### v2

@@ 4.4. Failure of the outgoing property: all fixed points on the axis of symmetry
From Theorem 4.3 we know that if a sufficiently smooth axisymmetric similarity profile $U\not\equiv 0$ solving (4.1) were to exist for some $\gamma<1/2$, then the local outgoing property of Definition 3.7 must fail. As discussed in Section 4.2 this failure can be either due to the fact that the nodal set $\mathcal{N}_{V}\subset\mathbb{R}^{3}$ is infinite (such a situation was discussed in Section 4.3), or because the local outgoing inequality (4.6) fails. In this section we consider the case that $\mathcal{N}_{V}\subset\mathbb{R}^{3}$ is finite, but the local outgoing inequality (4.6) does not hold. In particular, this implies that all fixed points of the ODE system (4.4) (the solutions of (4.7)) at which the swirl velocity vanishes, [^10] If the swirl velocity does not vanish at a solution of (4.7), Theorem 4.4 already gives $\gamma=1/2$. lie on the axis of symmetry $\{r=0\}$. [^11] For 3D incompressible Navier-Stokes, any putative finite time axisymmetric singularity must occur on the axis of symmetry [6]. Hence, if we are to lift axisymmetric self-similar blowups from incompressible 3D Euler to incompressible 3D Navier-Stokes, it is natural to search for Euler singularities which lie on $\vec{e}_{z}$. Our main result is:
@@
Let $U$ be a nontrivial $C^{2}$ smooth axisymmetric globally self-similar velocity profile for 3D incompressible Euler equations, solving (4.1), satisfying (3.5) and (3.8). Assume that all fixed points of the meridional Lagrangian flow (4.4) are isolated, and lie on the axis of symmetry $\{r=0\}$. Then $\gamma\geq 1/2$.
Notice that as opposed to Theorem 4.3, here we did not assume the validity of (4.6), nor did we assume the infinite smoothness of $\Omega$ at the nodal values.
@@
Assume by contradiction that $\gamma<1/2$. The fixed points of (4.4) on the axis of symmetry $\{r=0\}$ are precisely the zeros of the function $V_{z}(0,z)=\gamma z+U_{z}(0,z)$. Because $U(0)=0$ (cf. (3.5)), the origin is always a zero: $V_{z}(0,0)=0$. Furthermore, the sublinearity assumption (3.5) implies that $V_{z}(0,z)\sim\gamma z$ as $|z|\to\infty$, so $V_{z}(0,\cdot)$ has at most finitely many zeros.
Fix an arbitrary point $(r_{0},z_{0})$ with $r_{0}>0$, and consider its backward trajectory for the associated flow (4.4) in the meridional plane
$\Gamma_{0}=\{(R(\tau),Z(\tau))\colon\tau\leq 0\}=\{(R(r_{0},z_{0},\tau),Z(r_{0},z_{0},\tau))\colon\tau\leq 0\},$
@@
as $\tau\to$-$\infty$. We claim that $\Gamma_{0}$ is contained in a bounded set. Because $\mathcal{H}$ is non-increasing along forward trajectories (cf. (3.33)), backward in time $\mathcal{H}$ is non-decreasing. The far-field asymptotics (3.34) imply that the superlevel set $\mathcal{S}_{c}:=\{y\in\mathbb{R}^{3}:\mathcal{H}(y)\geq c\}$ is bounded for every $c\in\mathbb{R}$, since $\mathcal{H}(y)\to$-$\infty$ as $|y|\to\infty$ when $\gamma<1/2$. Setting $c=\mathcal{H}(r_{0},z_{0})$, the backward trajectory satisfies $\mathcal{H}(R(\tau),Z(\tau))\geq c$ for all $\tau\leq 0$, and is thus contained in $\Gamma_{0}\subset\mathcal{S}_{c}$.
@@
Next, we claim that the $\alpha$-limit set of $\Gamma_{0}$ is a single axis fixed point. The backward trajectory $\Gamma_{0}$ is bounded and lies in the two-dimensional meridional half-plane. By the Poincaré–Bendixson theorem, the $\alpha$-limit set $A$ is a nonempty, compact, connected, flow-invariant subset of the half-plane $\{r\geq 0,z\in\mathbb{R}\}$. As periodic orbits are excluded (on a periodic orbit $(\gamma r+U_{r},\gamma z+U_{z})$ is nonzero, so $\mathcal{H}$ would strictly decrease over one period, contradicting periodicity), $A$ must either be a single fixed point, or consist of multiple fixed points connected by heteroclinic/homoclinic orbits. Since all fixed points lie on the axis of symmetry $\{r=0\}$, any heteroclinic/homoclinic orbit in $A$ connecting axis fixed points would pass through $\{r>0\}$, violating Corollary 4.8. Therefore, $A=\{(0,z_{*})\}$ for some axis fixed point $z_{*}$, and the full trajectory converges: $(R(\tau),Z(\tau))\to(0,z_{*})$ as $\tau\to$-$\infty$.
Next, we note that the $U_{\theta}$ equation in (4.1) gives
$(\gamma r+U_{r})\partial_{r}(rU_{\theta})+(\gamma z+U_{z})\partial_{z}(rU_{\theta})+(1-2\gamma)(rU_{\theta})=0.$
Integrating this equation along the backwards trajectory $(R(\tau),Z(\tau))_{\tau\leq 0}$ results in
$e^{(1-2\gamma)\tau}R(\tau)U_{\theta}(R(\tau),Z(\tau))=r_{0}U_{\theta}(r_{0},z_{0}).$
Since $\gamma<1/2$, we have $e^{(1-2\gamma)\tau}\to 0$ as $\tau\to$-$\infty$. Previously we have shown $R(\tau)\to 0$ and $U_{\theta}(R(\tau),Z(\tau))\to U_{\theta}(0,z_{*})=0$ as $\tau\to$-$\infty$. Since $r_{0}>0$, we deduce $U_{\theta}(r_{0},z_{0})=0$; this fact is a manifestation of the conservation of circulation (3.27). With Remark 4.2, it follows that $U_{\theta}\equiv 0$, and hence also $\Omega_{r}\equiv\Omega_{z}\equiv 0$ on $0,\infty)\times\mathbb{R}$.
To conclude, we use the $\Omega_{\theta}$ equation in (4.3) and the fact that $U_{\theta}=0$ to deduce
$(\gamma r+U_{r})\partial_{r}\bigl(\tfrac{1}{r}\Omega_{\theta}\bigr)+(\gamma z+U_{z})\partial_{z}\bigl(\tfrac{1}{r}\Omega_{\theta}\bigr)+(1+\gamma)\bigl(\tfrac{1}{r}\Omega_{\theta}\bigr)=0.$
Integrating this equation along the backwards trajectory $(R(\tau),Z(\tau))_{\tau\leq 0}$ results in
$e^{(1+\gamma)\tau}\tfrac{1}{R(\tau)}\Omega_{\theta}(R(\tau),Z(\tau))=\tfrac{1}{r_{0}}\Omega_{\theta}(r_{0},z_{0}).$
We have $e^{(1+\gamma)\tau}\to 0$ as $\tau\to$-$\infty$. Moreover, as $\tau\to$-$\infty$ we have $R(\tau)\to 0$, and hence $\frac{1}{R(\tau)}\Omega_{\theta}(R(\tau),Z(\tau))\to(\partial_{r}\Omega_{\theta})(0,z_{*})$, which is finite by assumption ($\Omega$ is $C^{1}$). Thus, for any $r_{0}>0$ we deduce $\Omega_{\theta}(r_{0},z_{0})=0$. With Remark 4.2, it follows that $\Omega_{\theta}\equiv 0$ on $0,\infty)\times\mathbb{R}$. This contradicts the assumption that the solution $\Omega$ was nonzero; hence, we must have $\gamma\geq 1/2$. ∎